OpenStudy (anonymous):

The twice-differentiable function f is defined for all real numbers and satisfies the following conditions: f(0)=2, f'(0)=-4, and f''(0)=3

OpenStudy (anonymous):

a.) The function g is given by g(x)=e^(ax) + f(x) for all real numbers, where a is a constant. Find g'(0) and g''(0) in terms of a. Show the work that leads to your answers. b.) The function h is given by h(x)-cos(kx) f(x) for all real numbers, where k is a constant. Find h'(x) and write an equation for the line tangent to the graph of h at x=0

OpenStudy (anonymous):

\[g(x)=e^{ax}+f(x)\\ \implies g'(x)=ae^{ax}+f'(x)\\ g''(x)=a^2e^{ax}+f''(x)\]

OpenStudy (anonymous):

how would i find g'(0) and g''(0)

OpenStudy (sirm3d):

in \[g'(x)=ae^{ax}+f'(x)\] replace x by zero, and use the given information on f.

OpenStudy (sirm3d):

\[g'(0)\] means \[g'(x)\text{ when } x=0\]

OpenStudy (anonymous):

\[g'(0)=a \times e ^{a \times 0} - 4\]

OpenStudy (anonymous):

is that right?

OpenStudy (anonymous):

@agent0smith

OpenStudy (anonymous):

@RadEn

OpenStudy (raden):

yeah, u were right. just simplify that g′(0) = a * e^(a*0) − 4 g′(0) = a * e^0 − 4 g′(0) = a * 1 − 4 g′(0) = a − 4

OpenStudy (raden):

what about g"(0), did u get it ?

OpenStudy (anonymous):

g''(0)=a^2 + 3

OpenStudy (raden):

correct too

OpenStudy (anonymous):

how about for part b.)

OpenStudy (raden):

can u rewrite that function, please

OpenStudy (anonymous):

\[h(x)=\cos(kx)f(x)\]

OpenStudy (raden):

nah, that's good :)

OpenStudy (raden):

to get h' u have to use the product rule : if given h = f * g then h' = gf' +fg'

OpenStudy (anonymous):

is h'(x)=4cos(kx) + 4sin(kx)

OpenStudy (raden):

h(x)=cos(kx)f(x) so, h'(x) = f(x)(cos(kx))' + cos(kx)f '(x) h'(x) = k f(x)(-sin(kx)) + cos(kx)f '(x) h'(x) = -k f(x)sin(kx) + cos(kx)f '(x) i dont know why there is 4 like yours

OpenStudy (anonymous):

because the equation gives us f(0), f'(0), and f''(0)

OpenStudy (raden):

Oh question b still be related with a, sorry im forgot

OpenStudy (anonymous):

so was i correct? :P

OpenStudy (raden):

h'(x) = -k f(x)sin(kx) + cos(kx)f '(x) h'(0) = -k f(0)sin(k(0)) + cos(k(0))f '(0) h'(0) = -k (2)sin(0) + cos(0)f '(0) h'(0) = -k (2)(0) + 1(-4) h'(0) = 0 - 4 h'(0) = - 4

OpenStudy (raden):

woppps... missunderstanding :)

OpenStudy (raden):

hup, canceled.. i have right we just have h'(x) = -k f(x)sin(kx) + cos(kx)f '(x) only then to get slope (m), we have to find the value of h'(0) and i have done it above :P

OpenStudy (anonymous):

so f'(0)=-4 is correct?

OpenStudy (raden):

yes, h'(0) = -4 not f'(0)=-4

OpenStudy (anonymous):

so how do i write an equation for the line tangent to the graph of h at x=0

OpenStudy (raden):

there area 2 element to forms an equation of tangennt line are the slope (we got) and the other is a point slope of that curve in this case given a point slope x=0 (it just absis) then we have to determinen its ordinat point is y

OpenStudy (raden):

to get the value of y, just subtitute x=0 to h(x), in other words find h(0)

OpenStudy (raden):

h(x)=cos(kx)f(x) y = h(0) = cos(k(0))f(0) = cos(0)f(0) = 1 * 2 = 2

OpenStudy (raden):

so, the point slope at point (0,2)

OpenStudy (anonymous):

y=-4x+4

OpenStudy (raden):

now, use the formula straigh line equation :  with m is slope and (x1,y1) is the point slope